OpenStudy (anonymous):

dede and mike both run the length of a two-mile field. if dede runs 5 mph and mike runs 6 mph, how many more minutes does it take dede to run the field

OpenStudy (anonymous):

how long does it take to run 2 miles at 5 mph?

OpenStudy (anonymous):

For me it takes about 10 minutes to run 2 miles

OpenStudy (anonymous):

@satellite73: Have you seen that video on YouTube?

OpenStudy (anonymous):

if it is not clear, use "distance is rate times time" or \(D=RT\) and so time is distance divided by rate, aka \(T=\frac{D}{R}\)

OpenStudy (anonymous):

@FutureMathProfessor what is "that video"?

OpenStudy (anonymous):

You know which one I'm talking about! That guy quizzing his girlfriend about speeds in a car :)

OpenStudy (anonymous):

it would take me all morning to run two miles no, i haven't seen it

OpenStudy (anonymous):

http://www.youtube.com/watch?v=Qhm7-LEBznk

OpenStudy (anonymous):

can u explain how to get the answer

OpenStudy (anonymous):

what is t

OpenStudy (anonymous):

2 miles at five mile per hour takes \(\frac{2}{5}\) of an hour

OpenStudy (anonymous):

2 miles at 6 mpt takes \(\frac{2}{3}=\frac{1}{3}\) of an hour

OpenStudy (anonymous):

i know that but how to get the answer

OpenStudy (anonymous):

your job is to compute \[\frac{1}{3}-\frac{2}{5}\]

OpenStudy (anonymous):

that is how much longer it would take

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

is it 5/15-6/15

OpenStudy (anonymous):

actually that makes it more difficult, sorry compute \[\frac{2}{5}-\frac{1}{3}\]

OpenStudy (anonymous):

is 1/15

OpenStudy (anonymous):

got it one fifteenth of an hour now you have to convert to minutes: multiply by 60

OpenStudy (anonymous):

4 minutes

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw

OpenStudy (raden):

D1 = D2 R1T1 = R2T2 5T1 = 6T2 T1 = (6/5) T2 = (1 + 1/5)h = 1h + 1/5 * 60 m so, it takes about 12 minutes CMIIW

OpenStudy (anonymous):

so who is right

OpenStudy (radar):

Who do you think is right? Now that you know how to work it..........

OpenStudy (radar):

I think @satellite73 explanation was very reasonable. 2/5 of an hour is 24 minutes 2/6 of an hour is 20 minutes Whats the difference?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

whatever

OpenStudy (anonymous):

but thx for the help